A capacitor is charged to store an energy U. The charging battery is disconnected. An identical capacitor is now connected to the first capacitor in parallel. What is the energy in each capacitor

Grade:12

1 Answers

Khimraj
5 years ago
Let inItial voltage across capacitor is V. THen
½CV2 = U
and then final voltge across each capacitor is V1
Then
CV = CV1 + CV1
So V1 = V/2
Final energy in each capacitor = ½C(V/2)2 = ¼CV2.
Hope it clears. If you like answer then please approve it.